# One Model, Any Conjunctive Query: Graph Neural Networks for Answering Queries over Incomplete Knowledge Graphs

### v2

@@ 4.2 Query Answer Classification & Query Answer Retrieval
5 AnyCQ: framework for query answering
To address the introduced tasks of query answer classification and retrieval, we propose a neuro-symbolic framework for scoring arbitrary existential Boolean formulas, called AnyCQ. Let $\pi$ be a link predictor for an observable knowledge graph $G$. An AnyCQ model $\Theta$ equipped with $\pi$ can be viewed as a function $\Theta(G,\pi):\mathsf{CQ^{0}}(G)\rightarrow[0,1]$ where $\mathsf{CQ^{0}}(G)$ is the class of conjunctive Boolean queries over the same vocabulary as $G$. For input $Q=\exists\vec{y}.\Phi(\vec{y})$, $\Theta$ searches over the space of assignments to $\vec{y}$ for:
and returns an approximation $\Theta(Q|G,\pi)$ of $S_{\pi,G}(Q)$:
Note that unfolding the Boolean formula score gives:
$S_{\pi,G}(Q)=S_{\pi,G}(\Phi(\alpha_{\max}(\vec{y})/\vec{y}))\approx\Theta(Q|G,\pi).$
Hence, leveraging the strength of GNNs to find optimal solutions to combinatorial optimization problems, we can recover promising potential candidates for $\alpha_{\max}$, allowing to accurately estimate $S_{\pi,G}(Q)$.
Overview. During the search, our method encodes the query $Q$ and its relation to the current assignment $\alpha$ into a computational graph $G_{Q,\alpha}$ (Section˜5.1). This graph is then processed with a simple GNN $\theta$ (whose architecture is described in details in Section˜A.1), which updates its hidden embeddings and generates distributions $\mu$ from which the next assignment $\alpha^{\prime}$ is sampled (Section˜5.2).
5.1 Query representation
We transform the input queries into an undirected computational graph (Figure˜2), whose structure is adopted from ANYCSP (Tönshoff et al., 2023). Consider a conjunctive Boolean query $Q=\exists\vec{y}.\Phi(\vec{y})$ over a knowledge graph $G$, with $\Phi$ quantifier-free, and let $\pi$ be a link predictor for $G$. Let $c_{1},...,c_{n}$ be constant symbols mentioned in $\Phi$, and $\psi_{1},...,\psi_{l}$ be the literals in $\Phi$. We define the domain ${\mathcal{D}}(e)$ of the term $e$ as ${\mathcal{D}}(y)=V(G)$ for each existentially quantified variable $y$ and ${\mathcal{D}}(c_{i})=\{c_{i}\}$ for each constant $c_{i}$. Given an assignment $\alpha:\vec{y}\rightarrow\vec{a}$, the computational graph $G_{Q,\alpha}$ is constructed as follows:
@@ 5.1 Query representation
In contrast to PE labels, which are independent of the assignment $\alpha$, light edge (LE) labels reflect how local changes to $\alpha$ affect satisfiability of the literals. Formally, we set $L_{E}(v_{\psi_{i}},v_{e\rightarrow a};\alpha)=1$ if $\psi_{i}$ is satisfied under the assignment $\alpha_{z\rightarrow a}$, and 0 otherwise. In other words, LE labels answer the question: “If we change $\alpha$ so that $z$ is assigned to $a$, will $\psi_{i}$ be satisfied?”. Satisfiability is again determined by binarizing the prediction score returned by the link predictor $\pi$. Hence, through these edge labels, $\pi$ effectively guides the search toward promising updates in the assignment space. Further explanation of edge labels is provided in Section˜A.7.
@@
5.2 AnyCQ search process
The outline of the search process conducted by AnyCQ is presented in Figure˜3. Before the search commences, the hidden embeddings $\mathbf{h}^{(0)}$ of all value nodes are set to a pre-trained vector $\mathbf{h}\in\mathbb{R}^{d}$ and an initial assignment $\alpha^{(0)}$ is drafted, independently sampling the value for each variable $y\in\{\vec{y}\}$ uniformly at random from $\mathcal{D}(y)$. The variable and literal nodes are not assigned any hidden embeddings, serving as intermediate steps for value node embedding updates. At the beginning of search step $t$, $G_{Q,\alpha^{(t-1)}}$ is processed with a GNN $\theta$, which generates new value node embeddings $\mathbf{h}^{(t)}$, and for each variable $y\in\vec{y}$ returns a distribution $\mu_{y}^{(t)}$ over $\mathcal{D}(y)$. Finally, the next assignment $\alpha^{(t)}$ is sampled by drawing the value $\alpha^{(t)}(y)$ from $\mu_{y}^{(t)}$, independently for each $y\in\{\vec{y}\}$. A precise description of the architecture of $\theta$ is provided in Appendix A.1. The search terminates after $T$ steps, and the generated assignments $\alpha^{(0)},\alpha^{(1)},...,\alpha^{(T)}$ are used to approximate $S_{\pi,G}(Q)$:
5.3 Training
During training on each dataset, we equip AnyCQ with a predictor $\pi_{\text{train}}$, representing the training graph $G_{\text{train}}$. Thus, the only trainable component of $\Theta$ remains the GNN $\theta$. We utilize the training splits from the existing CQA datasets Ren and Leskovec (2020), hence limiting the scope of queries viewed during training to formulas mentioning at most three variables. Moreover, we restrict the number of search steps $T$ to at most $15$, encouraging the network to quickly learn to apply logical principles locally. Inspired by prior work on combinatorial optimization (Shi and Zhang, 2022; Tönshoff et al., 2023; Abe et al., 2019), we train $\theta$ in a reinforcement learning setting via REINFORCE Williams (2004), treating $\theta$ as a search policy network with the objective of maximizing $\Theta(Q|G,\pi)$. This setup enables AnyCQ to generalize across different query types, scaling to formulas of size several times larger than observed during training, as demonstrated in Section˜6. The complete methodology is presented in Section˜A.2.
@@ 5.4 Theoretical and conceptual properties
AnyCQ is complete given sufficiently many search steps – any AnyCQ model equipped with any link predictor will eventually return the corresponding Boolean formula score:
Theorem 5.1.
(Completeness) Let $Q=\exists\vec{y}.\Phi(\vec{y})$ be a conjunctive Boolean query and let $\Theta$ be any AnyCQ model equipped with a predictor $\pi$. For any execution of $\Theta$ on $Q$, running for $T$ steps:
When AnyCQ is equipped with the perfect link predictor for $\tilde{G}$, we can guarantee the soundness of predictions, i.e., all positive answers will be correct (proofs in Appendix˜B):
Theorem 5.2.
(Soundness) Let $Q\!=\!\exists\vec{y}.\Phi(\vec{y})$ be a conjunctive Boolean query over a knowledge graph $G$ and let $\Theta$ be any AnyCQ model equipped with a perfect link predictor $\tilde{\pi}$ for $\tilde{G}$. If $\Theta(Q|G,\tilde{\pi})\!>\!0.5$, then $\tilde{G}\!\models\!Q$.
Transferability. By construction, our trained GNN $\theta$ is independent of the equipped link predictor $\pi$ and the knowledge graph $G$. Therefore, once trained, an AnyCQ model can be applied to answer queries over any dataset, with the only necessary augmentation being the change of the equipped link predictor. We validate this transferability in Section˜6.4, comparing the performance of models trained on FB15k-237 and NELL datasets on the QAR task.
Generality. Although this work focuses on knowledge graphs with binary relations, by construction, AnyCQ can handle predicates with arbitrary arity requiring no changes in its structure. Assuming the availability of a relevant link predictor, AnyCQ can hence be applied to hyper-relational data. Similarly, by equipping a fully inductive link predictor (Galkin et al., 2024a), AnyCQ can serve as a general query answering engine without the need for interchanging the equipped predictor between input samples. Experiments with equipping Ultra (Section˜D.5) already point out the potential of AnyCQ in the inductive setting. Importantly, our method is not limited to scoring only conjunctive Boolean queries and can process all formulas in conjunctive or disjunctive normal form (see Section˜A.4).
@@ 6 Experimental evaluation
We empirically evaluate AnyCQ on the tasks of QAC (Section˜6.2) and QAR (Section˜6.3). We also conduct two ablation studies (Section˜6.4): first, we assess AnyCQ with a perfect link predictor to isolate and measure the quality of the search engine independent of the predictor’s imperfections; second, we examine the generalizability of AnyCQ by applying it to out-of-distribution KGs.
Figure 4: Examples of query graphs of formulas from our FB15k-237-QAR benchmark. Blue nodes represent constants, grey nodes - distinct existentially quantifies variables, and orange nodes - free variables.
@@ 6.1 Experimental setup
Benchmarks and datasets. Existing benchmarks (Ren and Leskovec, 2020; Yin et al., 2023) comprise formulas with simple structures, thereby impeding the comprehensive evaluation and advancement of novel methodologies. We address this gap by creating new datasets on top of well-established benchmarks, consisting of queries with demanding structural complexity. Specifically, we generate formulas mentioning up to 20 distinct terms, allowing the presence of multiple cycles, non-anchored leaves, long-distance reasoning steps, and multi-way conjunctions (see Figure˜4). The generation process is described in Appendix˜C. Building on the existing CQA datasets, we propose two novel benchmarks for QAC: FB15k-237-QAC and NELL-QAC, each divided into 9 splits, consisting of small and large formulas. For the task of QAR, we observe that many instances of the simpler query structures, inherited from existing CQA benchmarks, admit easy answers, i.e. have at least one satisfying assignment supported entirely by observed facts. Combined with their limited structural complexity, this makes them trivial under the QAR objective, which only requires recovering a single correct answer. To evaluate reasoning under incompleteness and structural difficulty, we introduce new benchmarks: FB15k-237-QAR and NELL-QAR, consisting of large formulas with up to three free variables. In QAC, we focus exclusively on single-variable instances, as multi-variable cases reduce trivially to the single-variable setting. For example, $\langle Q(x_{1},x_{2}),(a_{1},a_{2}),G\rangle$ is equivalent to a single-variable instance $\langle Q(x_{1},a_{2}/x_{2}),a_{1},G\rangle$, as they both ask if $\tilde{G}\models Q(a_{1}/x_{1},a_{2}/x_{2})$.
@@
Baselines. As the baselines for the small-query split on our QAC task, we choose the state-of-the-art solutions from CQA capable of handling the classification objective: QTO (Bai et al., 2023), FIT (Yin et al., 2024), GNN-QE (Zhu et al., 2022) and UltraQuery (Galkin et al., 2024b). Considering the large-query splits, we were surprised to notice that no existing approaches can be applied in this setting, as none of them can simultaneously: 1) efficiently handle cyclic queries and 2) produce calibrated probability estimates, without the knowledge of the trivial answers.
Indeed, standard methods like BetaE (Ren and Leskovec, 2020), CQD (Arakelyan et al., 2020), ConE (Zhang et al., 2021), GNN-QE (Zhu et al., 2022) or QTO (Bai et al., 2023) are limited to tree-like queries. Neural approaches, such as LMPNN (Wang et al., 2022) or CLMPT (Zhang et al., 2024), are trained using Noisy Contrastive Estimation; hence their predictions do not meaningfully translate to desired probabilities. Finally, FIT (Yin et al., 2024) and Q2T (Xu et al., 2023) require transforming scores predicted by their ComplEx-based link predictors, while all known schemes (see Section˜D.1) assume the set of easy answers is known, or otherwise, trivial to recover. Hence, we furthermore use an SQL engine, implemented by DuckDB (Raasveldt and Mühleisen, 2019), reasoning over the observable graph. For the same reasons, extended by the need of reasoning over queries with multiple variables, we consider only the SQL engine as the baseline for QAR experiments. In both cases, we limit the processing time to 60 seconds, ensuring termination in a reasonable time. Additional evaluations ablating the impact of this timeout, using 30, 60, and 120 seconds thresholds, are included in Appendix˜E. Training details for the considered baselines are provided in Appendix˜D.
@@
Methodology. Given a Boolean query $Q$ over an observable KG $G$, an AnyCQ model $\Theta$ equipped with a link predictor $\pi$ for $G$ can decide if $\tilde{G}\models Q$, by returning whether $\Theta(Q|G,\pi)>0.5$. We use this functionality to solve QAC instances by applying our AnyCQ models directly to $Q(\vec{a}/\vec{x})$. For the QAR task, given a query $Q(\vec{x})$ over an observable KG $G$, we run our AnyCQ framework on the Boolean formula $\exists\vec{x}.Q(\vec{x})$, returning $\mathsf{None}$ if the returned $\Theta(\exists\vec{x}.Q(\vec{x})|G,\pi)$ was less than $0.5$. Otherwise, we return $\alpha(\vec{x})$ where $\alpha$ is the visited assignment maximizing the Boolean formula score. In both scenarios, we perform 200 search steps on each input instance in the large query splits, and just 20 steps for small QAC queries. We equip NBFNet as the link predictor for both QAC and QAR evaluations (details in Section˜D.5).
@@
Metrics. Given the classification nature of both our objectives, we use the F1-score as the metric for query answer classification and retrieval (see Section˜C.4 for details). In QAR, we mark a positive solution as correct only if the returned assignment is an answer to the input query. In contrast to the CQA evaluation, we do not distinguish between easy and hard answers, since the task of efficiently answering queries with advanced structural complexity, even admitting answers in the observable knowledge graph, is not trivial.
Table 1: Average F1-scores of considered methods on the query answer classification task.
@@
Table 2: F1-scores on the QAR datasets. $k$ is the number of free variables in the input query.
@@
6.2 Query answer classification experiments
@@
Table 3: Recall on the easy and hard samples from the QAR datasets. $k$ is the number of free variables.
Dataset Model 3-hub 4-hub 5-hub
Easy instances
FB15k-237-QAR SQL 88.7 61.2 26.4 62.8 87.2 71.7 52.6 71.9 85.3 63.6 61.4 70.5
AnyCQ 83.3 84.2 71.6 80.5 79.1 71.1 72.1 74.3 74.0 67.0 52.3 65.0
NELL-QAR SQL 94.5 69.0 55.4 79.6 90.1 69.0 63.4 77.9 88.8 78.6 64.0 80.2
AnyCQ 93.0 89.4 86.5 90.7 89.6 88.8 79.6 87.1 89.4 84.5 81.3 86.1
Hard instances
FB15k-237-QAR SQL 0 0 0 0 0 0 0 0 0 0 0 0
AnyCQ 11.7 7.8 11.1 10.2 8.8 7.0 6.5 7.4 16.8 10.8 8.0 11.7
NELL-QAR SQL 0 0 0 0 0 0 0 0 0 0 0 0
AnyCQ 7.0 7.0 4.0 5.9 9.7 9.2 8.2 9.0 9.2 8.1 5.8 7.7
6.3 Query answer retrieval experiments
We present the QAR evaluation results across all splits of the two proposed datasets consisting of large formulas with multiple free variables in Table˜2. Compared with the SQL engine which can only extract easy answers, AnyCQ can reliably achieve similar performance on easy answers while also extracting hard answers over these large formulas, thus outperforming the standard SQL engine in almost all metrics. Importantly, as the number of free variables k increases, the performance gap between AnyCQ and SQL becomes more pronounced. This improvement is due to AnyCQ’s computation graphs that effectively handle the rising query complexity without sacrificing performance, substantially gaining over traditional solvers.
Easy and hard recall on QAR. To further interpret the results on our query answer retrieval benchmarks, we analyze the recall metric, distinguishing between easy and hard samples (Table˜3). As anticipated, the performance of the SQL engine degrades as the number of free variables in the input query increases. While a similar behavior can be witnessed for AnyCQ models, it progresses at a much slower rate, and as a consequence, AnyCQ consistently outperforms SQL on nearly all splits involving more than one free variable. Additionally, for unary queries, AnyCQ’s recall remains within $15\%$ relative to SQL, retrieving answers for over $70\%$ of trivial queries across most splits.
As shown in Table˜3, AnyCQ is capable of finding non-trivial answers, even to complicated queries. In contrast, the SQL engine, as a classical approach, cannot retrieve answers to queries that lack a correct solution in the observable knowledge graph. AnyCQ, leveraging a link predictor, demonstrates the ability to retrieve unobserved yet correct answers, even for large, structurally complex queries with multiple free variables.

### v3

@@ 5.1 Query representation
In contrast to PE labels, which are independent of the assignment $\alpha$, light edge (LE) labels reflect how local changes to $\alpha$ affect satisfiability of the literals. Formally, we set $L_{E}(v_{\psi_{i}},v_{e\rightarrow a};\alpha)=1$ if $\psi_{i}$ is satisfied under the assignment $\alpha_{z\rightarrow a}$, and 0 otherwise. In other words, LE labels answer the question: “If we change $\alpha$ so that $z$ is assigned to $a$, will $\psi_{i}$ be satisfied?”. Satisfiability is again determined by binarizing the prediction score returned by the link predictor $\pi$. Hence, through these edge labels, $\pi$ effectively guides the search toward promising updates in the assignment space.
Further explanations of edge labels are provided in Section˜B.8.
5.2 AnyCQ search process
The outline of the search process conducted by AnyCQ is presented in Figure˜3. Before the search commences, the hidden embeddings $\mathbf{h}^{(0)}$ of all value nodes are set to a pre-trained vector $\mathbf{h}\in\mathbb{R}^{d}$ and an initial assignment $\alpha^{(0)}$ is drafted, sampling the value for each variable $y\in\{\vec{y}\}$ independently and uniformly at random from $\mathcal{D}(y)$. The variable and literal nodes are not assigned any hidden embeddings, serving as intermediate steps for value node embedding updates. At the beginning of search step $t$, $G_{Q,\alpha^{(t-1)}}$ is processed with a GNN $\theta$, which generates new value node embeddings $\mathbf{h}^{(t)}$, and for each variable $y\in\vec{y}$ returns a distribution $\mu_{y}^{(t)}$ over $\mathcal{D}(y)$. Finally, the next assignment $\alpha^{(t)}$ is sampled by drawing the value $\alpha^{(t)}(y)$ from $\mu_{y}^{(t)}$, independently for each $y\in\{\vec{y}\}$. A precise description of the architecture of $\theta$ is provided in Appendix B.1. The search terminates after $T$ steps, producing assignments $\alpha^{(0)},\alpha^{(1)},...,\alpha^{(T)}$ which are used to approximate $S_{\pi,G}(Q)$ as
5.3 Training
During training on each dataset, we equip AnyCQ with a predictor $\pi_{\text{train}}$, representing the training graph $G_{\text{train}}$. Thus, the only trainable component of $\Theta$ remains the GNN $\theta$. We utilize the training splits from the existing CQA datasets [betae], hence limiting the scope of queries viewed during training to formulas mentioning at most three variables. Moreover, we restrict the number of search steps $T$ to at most $15$, encouraging the network to quickly learn to apply logical principles locally. Inspired by prior work on combinatorial optimization [neuraltsp, anycsp, Abe2019SolvingNP], we train $\theta$ in a reinforcement learning setting via REINFORCE [reinforce], treating $\theta$ as a search policy network with the objective of maximizing $\Theta(Q|G,\pi)$. This setup enables AnyCQ to generalize across different query types, scaling to formulas of size several times larger than observed during training. See complete methodology in Section˜B.2.
@@ 5.4 Theoretical and conceptual properties
Our AnyCQ framework is supported by strong theoretical guarantees and designed for broad conceptual flexibility. Theoretically, the method is provably complete, ensuring it converges to the correct answer with sufficient search steps (Theorem˜C.1), and it is sound when equipped with a perfect link predictor, guaranteeing the correctness of its positive predictions (Theorem˜C.3). Conceptually, AnyCQ is built for transferability and generality. Because its core model is independent of the input graph and link predictor, it can be seamlessly applied to unseen knowledge graphs, as we have shown in Section˜6.3. Furthermore, its general design allows it to handle relations of any arity and process complex formulas in conjunctive or disjunctive normal form (Section˜B.5), even extending to inductive settings with the appropriate predictor (Section˜E.5).
@@ 6 Experimental evaluation
We empirically evaluate AnyCQ to assess its performance on the proposed tasks of Query Answer Classification (QAC) and Query Answer Retrieval (QAR). To provide a comprehensive analysis, we aim to answer the following key questions:
  • Q1.
    How does AnyCQ perform on QAC and QAR comparing with existing models? (Sections˜6.2 and 6.2)
  • Q2.
    How does AnyCQ perform outside the training domain? (Section˜6.3)
  • Q3.
    How does AnyCQ perform with a perfect link predictor? (Section˜6.3)
  • Q4.
    How does the choice of link predictor affect AnyCQ’s overall performance? (Appendix˜E)
  • Q5.
    How does AnyCQ’s performance scale with increasing query complexity? (Appendix˜F)
@@ 6.1 Experimental setup
Benchmarks and datasets. Existing CQA benchmarks [betae, efok_dataset] contain mostly simple query structures, which impair development of more advanced and general methods. To close this gap, we generate new datasets on top of standard benchmarks, introducing queries of higher structural complexity. These formulas may involve up to 20 distinct terms and feature multiple cycles, non-anchored leaves, long-range dependencies, and multi-way conjunctions. See Appendix˜D for generation details.
For QAC, we propose FB15k-237-QAC and NELL-QAC, each divided into 9 splits, consisting of small and large formulas. We focus exclusively on single-variable instances, as multi-variable cases reduce trivially to the single-variable setting, i.e., $\langle Q(x_{1},x_{2}),(a_{1},a_{2}),G\rangle$ is equivalent to a single-variable instance $\langle Q(x_{1},a_{2}/x_{2}),a_{1},G\rangle$ as they both ask if $\tilde{G}\models Q(a_{1}/x_{1},a_{2}/x_{2})$.
@@
For QAR, we observe that many instances of the simpler query structures, inherited from existing CQA benchmarks, admit easy answers, i.e. have at least one satisfying assignment supported entirely by observed facts. Combined with their limited structural complexity, this makes them trivial under the QAR objective, which only requires recovering a single correct answer. To evaluate reasoning under incompleteness and structural difficulty, we introduce new benchmarks: FB15k-237-QAR and NELL-QAR, consisting of large formulas with up to three free variables.
@@
Baselines. As the baselines for the small-query split on our QAC task, we choose the state-of-the-art solutions from CQA capable of handling the classification objective: QTO [qto], FIT [fit], GNN-QE [gnn-qe] and UltraQuery [galkin2024ultraquery]. Considering the large-query splits, we notice that no existing approaches can be applied in this setting, as none of them can simultaneously: 1) efficiently handle cyclic queries and 2) produce calibrated probability estimates, without the knowledge of the trivial answers [^2] See Section A.2 for a detailed explanation.. Hence, we furthermore use an SQL engine, implemented by DuckDB [duckdb], reasoning over the observable graph. For the same reasons, extended by the need of reasoning over queries with multiple variables, we consider only the SQL engine as the baseline for QAR experiments. In both cases, we limit the processing time to 60 seconds, ensuring termination in a reasonable time. Additional evaluations ablating the impact of this timeout, using 30, 60, and 120 seconds thresholds, are included in Appendix˜F. Training details for the considered baselines are provided in Appendix˜E.
@@
Methodology. Given a Boolean query $Q$ over an observable KG $G$, an AnyCQ model $\Theta$ equipped with a link predictor $\pi$ for $G$ can decide if $\tilde{G}\models Q$, by returning whether $\Theta(Q|G,\pi)>0.5$. We use this functionality to solve QAC instances by applying our AnyCQ models directly to $Q(\vec{a}/\vec{x})$. For the QAR task, given a query $Q(\vec{x})$ over an observable KG $G$, we run our AnyCQ framework on the Boolean formula $\exists\vec{x}.Q(\vec{x})$, returning $\mathsf{None}$ if the returned $\Theta(\exists\vec{x}.Q(\vec{x})|G,\pi)$ was less than $0.5$. Otherwise, we return $\alpha(\vec{x})$ where $\alpha$ is the visited assignment maximizing the Boolean formula score. In both scenarios, we perform 200 search steps on each input instance in the large query splits, and just 20 steps for small QAC queries. We equip NBFNet as the default link predictor for both QAC and QAR evaluations (details in Section˜E.5).
@@
Metrics. Given the classification nature of both our objectives, we use the F1-score as the metric for query answer classification and retrieval (see Section˜D.4 for details). In QAR, we mark a positive solution as correct only if the returned assignment is an answer to the input query. In contrast to the CQA evaluation, we also include easy answer (instances), since the task of efficiently answering queries with advanced structural complexity, even admitting answers in the observable knowledge graph, is not trivial. The code can be found in this GitHub repository.
Table 1: Average F1-scores of considered methods on the QAC task.
@@
Table 2: F1-scores on QAR samples, where $k$ is the number of free variables.
@@
6.2 Main experiments results over QAC and QAR
Query answer classification (QAC) experiments
@@ 6.2 Main experiments results over QAC and QAR
Query answer retrieval (QAR) experiments
We present the QAR evaluation results across all splits of the two proposed datasets consisting of large formulas with multiple free variables in Table˜2. The performance of the SQL engine degrades as the number of free variables in the input query increases. While a similar behavior can be witnessed for AnyCQ models, it progresses at a much slower rate. Furthermore, AnyCQ is capable of finding non-trivial answers, even to complicated queries. As a consequence, AnyCQ consistently outperforms SQL on all splits, with the biggest differences being witnessed for queries involving more than one free variable.
A further analysis (detailed in Appendix˜F) shows that AnyCQ does not fall behind SQL on instances admitting observable answers, remaining within $10\%$ relative to the classical engine on unary queries, while outperforming it on multivariate splits. Moreover, AnyCQ correctly solves a fair share of hard instances, demonstrating its ability to retrieve unobserved yet correct answers, even for large, structurally complex queries with multiple free variables.